OpenStudy (anonymous):

Factor completely 18x^2 - 21x – 15. 3(2x + 1)(3x - 5) 3(2x - 5)(3x + 1) 3(2x - 1)(3x + 5) 3(6x + 1)(x - 5)

OpenStudy (nurali):

3(6x² - 7x - 5) Then, by factoring decomposition: 3(6x² + 3x - 10x - 5) = 3(3x(2x + 1) - 5(2x + 1)) = 3(2x + 1)(3x - 5)